1. A bunch of AMDM students were able to squeeze 50 people in a 9 x 9 ft square. Using this ratio, how much space does one person use?

2. Using this ratio, how many people can fit in a hallway that measures 16 x 75 feet?

Answer by stanbon(75887)

1. A bunch of AMDM students were able to squeeze 50 people in a 9 x 9 ft square. Using this ratio, how much space does one person use?
----
space per person = (9 x 9)/50 = 1.62 sq. ft
----------------------------------------------
2. Using this ratio, how many people can fit in a hallway that measures 16 x 75 feet?
# = (16 * 75)/1.62 = 741 when rounded up.
--------------
Cheers,
Stan H.
-----------------------